# Beyond the Visible: A Survey on Cross-spectral Face Recognition

### v3

@@ 10 Conclusions
In this article, we reviewed methods that have been developed for cross-spectral face recognition (CFR), placing emphasis on recent deep CNN-architectures. We discussed associated benefits and limitations, as well as the employed datasets. CFR is endowed with the ability to identify people in challenging illumination settings, as well as to detect presentation attacks. Such benefits have been exploited in applications such as defense, surveillance and public safety.
@@
While traditional FR is restricted to compare images in the visible spectrum, CFR is inherently more challenging, as both subject identification as well as spectral variation have to be addressed. In order to separate spectral variation from identity, approaches such as domain-invariant feature representation have been proposed. Despite such CNN-based approaches usually performing well in minimizing spectral variation between NIR and visible spectral bands, the large modality gap introduced by capturing face images in the thermal spectral band can be overcome employing GANs.
In CNN-based approaches, streamlined to learn domain-invariant features, the triplet loss function has demonstrated better performance than the contrastive loss function, since the former considers both intra-class minimization and inter-class maximization. For GAN-based approaches, the identity loss function is essential to ensure that synthesized faces preserve the same identity as the ground-truth face. Other loss functions such as attribution loss, shape loss or semantic loss are utilized as auxiliary loss functions.
@@
In case of spectral band selections, NIR is preferred for reflective IR-to-Visible FR in the majority of applications. There is not much difference between MWIR or LWIR spectral bands in emissive IR-to-Visible face recognition. Nevertheless, the use of polarimetry for LWIR has been attributed to better face comparison results. While NIR and SWIR spectral bands can produce relatively good face resolutions, this remains a challenge for MWIR and LWIR sensors. Towards improving the poor resolution associated with thermal images, there is a necessity to develop high-resolution sensors by integrating super-resolution post-processing schemes. CFR has been impeded by lack of large publicly available face datasets. With the emergence of more comprehensive and large-scale datasets, such as ARL-VTF [111] and MILAB-VTF(B) [106], we foresee the performance of CFR further improving.
@@
To conclude, CFR has several challenging issues, and this survey has shown that the research community has an urgent need for systematic theoretical research in that field, rather than empirical experimentation with different loss functions and networks. Such concerns necessitate the introduction of robust and reliable methods to identify forged images and videos.

### v4

@@ 9 Conclusions
In this survey, we reviewed methods that have been developed for cross-spectral face recognition (CFR), placing emphasis on recently proposed deep CNN-architectures. We discussed associated benefits and limitations, as well as the datasets used. CFR is endowed with the ability to identify people in challenging illumination settings, as well as to detect presentation attacks. Such benefits have been exploited in applications such as defense, surveillance and public safety.
@@
While traditional FR is restricted to comparing images in the visible spectrum, CFR is inherently more challenging, as it must address both subject identification and spectral variation. In order to separate spectral variation from identity, approaches such as domain-invariant feature representation have been proposed. Despite such CNN-based approaches usually performing well in minimizing spectral variation between NIR and visible spectral bands, the large modality gap introduced by capturing face images in the thermal spectral band can be effectively addressed using GANs. In CNN-based approaches designed to learn domain-invariant features, the triplet loss function has demonstrated better performance than the contrastive loss function, since the former considers both intra-class minimization and inter-class maximization. For GAN-based approaches, the identity loss function is essential to ensure that synthesized faces preserve the same identity as the ground-truth face. Other loss functions such as attribution loss, shape loss, perceptual loss or semantic loss are utilized as auxiliary loss functions.
We here conclude that the NIR-to-visible CFR scenario has been extensively studied in the literature, achieving high accuracy. However, SWIR and MWIR sub-bands remain widely unused due to their limited practicality and high equipment costs. We note that the COVID-19 pandemic has brought to the fore a significant decrease in the cost of LWIR sensors, rendering such sensors more accessible for a wider range of applications. This, coupled with the increasing demand for surveillance and protection solutions, e.g., against spoofing attacks, has made LWIR imaging an appealing option for various industries.
@@
In terms of spectral band selection, NIR is preferred for reflective IR-to-Visible FR in the majority of applications. There is not much difference between MWIR or LWIR spectral bands in emissive IR-to-Visible face recognition. Nevertheless, the use of polarimetry for LWIR has been attributed to better face comparison results. While NIR and SWIR spectral bands can produce relatively good face resolutions, this remains a challenge for MWIR and LWIR sensors. Towards improving the poor resolution associated with thermal images, there is a necessity to develop high-resolution sensors by integrating super-resolution post-processing schemes. CFR has been impeded by lack of large publicly available face datasets. With the emergence of more comprehensive and large-scale datasets, such as ARL-VTF [113] and MILAB-VTF(B) [108], we foresee further improvements in CFR performance.
@@
To conclude, CFR has several challenging issues, and this survey has reiterated the significant need for systematic theoretical research in this field, rather than empirical experimentation with different loss functions and networks. Such concerns necessitate the introduction of robust and reliable methods to identify forged images and videos.